OpenStudy (anonymous):

solve the equation 3/4a=12

OpenStudy (anonymous):

solve for a a=12/(3/4) or a=12/0.75 That's if your equation was (3/4)a=12

OpenStudy (anonymous):

a = 12/3*4 a = 16

OpenStudy (anonymous):

Agree with bcurry333